OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). A:y = 2x + 6 B:y =−1/2x + 3 C:y =−1/2x +6 D: y = 2x + 3

OpenStudy (anonymous):

I really need help on this.. i have no idea how to do this.. Help! :(

OpenStudy (anonymous):

so to get the equation of a perpindicular line, you would have to flip the slope and make it negative which would get you -1/2 for your slope.

OpenStudy (anonymous):

then you would use the slope to find out your y intercept which will equal 6

OpenStudy (anonymous):

So the answer would be B or C?

OpenStudy (anonymous):

So C is the answer?

OpenStudy (anonymous):

exactly!!

OpenStudy (anonymous):

I got one more question if you don't mind?

OpenStudy (anonymous):

hit me!!

OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3).

OpenStudy (anonymous):

ok so when you have a line that is parrallel to the other the slope will stay the same!

OpenStudy (anonymous):

Hold on i have some answer choices

OpenStudy (anonymous):

A:y = 2x − 7 B: y =−1/2x − 2 C:y = 2x − 2 D:y =−1/2x -7

OpenStudy (anonymous):

it should be c!!

OpenStudy (anonymous):

Thank you <3

OpenStudy (anonymous):

you are very welcome!!